1. A bullet is fired from gun mounted at an angle of 30°. If the muzzle velocity is



400 m/s, calculate:



a. Vertical velocity component



b. Horizontal velocity component



c. Maximum height reach



d. Time of flight



e. Range



p.s.with calculation if possible please, thank you, sir:)

1
Expert's answer
2021-10-31T18:08:40-0400

"v_y=400\sin{30}-9.8t\\v_x=400\cos{30}=\frac{m}{s}\\H=\frac{400^2\sin^2{30}}{2(9.8)}=2041\ m\\\\t=2\frac{400\sin{30}}{(9.8)}=40.8\ s\\\\R=\frac{400^2\sin{60}}{(9.8)}=14139\ m\\"